# Low-Rank Quantization-Aware Training for LLMs

### v2

@@ Abstract
Large language models (LLMs) are omnipresent, however their practical deployment is challenging due to their ever increasing computational and memory demands. Quantization is one of the most effective ways to make them more compute and memory efficient. Quantization-aware training (QAT) methods, generally produce the best quantized performance, however it comes at the cost of potentially long training time and excessive memory usage, making it impractical when applying for LLMs. Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose LR-QAT – a lightweight and memory-efficient QAT algorithm for LLMs. LR-QAT employs several components to save memory without sacrificing predictive performance: (a) low-rank auxiliary weights that are aware of the quantization grid; (b) a downcasting operator using fixed-point or double-packed integers and (c) checkpointing. Unlike most related work, our method (i) is inference-efficient, leading to no additional overhead compared to traditional PTQ; (ii) can be seen as a general extended pretraining framework, meaning that the resulting model can still be utilized for any downstream task afterwards; (iii) can be applied across a wide range of quantization settings, such as different choices quantization granularity, activation quantization, and seamlessly combined with many PTQ techniques. We apply LR-QAT to LLaMA-2/3 and Mistral model families and validate its effectiveness on several downstream tasks. Our method outperforms common post-training quantization (PTQ) approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage. Specifically, we can train a 7B LLM on a single consumer grade GPU with 24GB of memory.
@@ 1 Introduction
Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose Low-Rank Quantization-Aware Training (LR-QAT) – a lightweight memory-efficient and inference-efficient QAT algorithm for LLMs. LR-QAT reduces the memory requirements of training a 7B LLM from >70GB of GPU memory to <21GB without degrading the predictive performance compared to traditional full-model QAT, making it possible to train such models on a single consumer grade GPU. Unlike most related work that combines low-rank adaptation with quantization, our method is also inference-efficient. After the training is complete, the auxiliary matrices are naturally absorbed into the quantized weight tensor without loss of accuracy and no extra overhead at inference time. LR-QAT is positioned as a general extended pretraining method, as opposed to being strictly a fine-tuning method – the resulting model is a low-bit general pretrained LLM, that can still be utilized for any task afterwards. If needed, our resulting low-bit pretrained LLM can be fine-tuned on specific downstream tasks or used with multiple LoRA adapters for rapid switching between various tasks.
@@
We apply LR-QAT to LLaMA-2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks. Our method outperforms recent PTQ approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage. Finally, our method can be applied across a wide range of quantization settings, including per-channel or per-block weight quantization, activation quantization, and can be combined with most of other PTQ techniques.
@@ 2 Background and related work
Figure 1: Left: A schematic illustration of our proposed LR-QAT. $\mathbf{x}$ and $\mathbf{y}$ denote the input and the output of the linear layer. Right: Memory requirements for training with various QAT techniques on LLaMA-2 7B (assuming batch size 1, sequence length 1024, $r=32$, and BF16 compute data type).
@@ 3 Method
Let’s recall how traditional QAT works. Given a linear layer with a weight matrix $\bm{W}\in\mathbb{R}^{m\times k}$ and assuming $b$-bit symmetric uniform affine quantization, the quantization is simulated as follows:
@@
Note that both in the original LoRA paper [25] and in the related work like QLoRA [16], there is no need to compute the product $\bm{A}\bm{B}$. Instead, those matrices are multiplied with the activations $\mathbf{x}$ as $\bm{A}\left(\bm{B}\mathbf{x}\right)$. However, we do compute a product $\bm{A}\bm{B}$ in (5), and in the naïve implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage. To prevent this, we employ gradient checkpointing [10] on (5). In other words, we recompute the quantizer function in the backward pass, leading to a slight runtime overhead but avoiding significantly increased memory usage.
@@ 4 Experiments
$\varphi\!\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. W4pc W3pc
WT-2 $\downarrow$ zero-shot $\uparrow$ WT-2 $\downarrow$ zero-shot $\uparrow$
FP32 FP32 LoRA 5.69 69.28 6.21 66.62
FP16 FP32 LoRA $+$0.00 --0.13 $+$0.01 --0.01
BF16 FP32 LoRA --0.01 $+$0.11 $+$0.01 $+$0.45
Q4.4 / Q3.5 FP32 LoRA --0.01 $+$0.16 $+$0.01 $+$0.31
Q4.4 / Q3.5 BF16 LoRA --0.01 $+$0.15 $+$0.01 $+$0.31
INT-4 / INT-3 FP32 LoRA $+$0.02 --0.04 $+$1.5e3 --30.30
INT-4 / INT-3 FP32 LoftQ ($T=1$) $+$0.14 --1.02 $+$0.18 $+$0.26
INT-4 / INT-3 FP32 LoftQ ($T=64$) $+$0.40 --1.40 $+$1.37 --2.01
@@
We assess the effectiveness of LR-QAT by conducting experiments on LLaMA-2 7B/13B [61], LLaMA-3 8B [1] and Mistral-0.1 7B [31]. To compare to relevant literature with weight-activation quantization, we further experiment with LLaMa-1 7B [60]. We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\!\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$. We then compare our method in terms of accuracy to standard full-model QAT, other baselines, and the related work. All detailed hyperparameters of our experiments are in Appendix B.
Quantization
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We quantize all linear layers, except the classification head. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [55], we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
Evaluation
Following the previous work [18, 65, 55, 44], we evaluate quantized models by reporting the perplexity of language generation on WikiText-2 [48], using a sequence length of $2048$. We also report zero-shot accuracy on a set of common sense reasoning tasks including BoolQ [12], PIQA [5], Winogrande [54], ARC [13], and HellaSwag [69]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
Datasets and training
We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network). We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$. We train on a small subset of SlimPajama [57], which is an open-source dataset similar to the original one used for pretraining LLaMA models. In all experiments we train using batch size $32$ and a maximum sequence length of $1024$. For weight-only and weight-activation quantization results, we train for $10^{4}$ steps. For ablation studies, we use shorter training of $10^{3}$ steps. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
Baselines
We compare with round-to-nearest quantization (RTN), where we set the ranges based on minimizing the $L^{p}$-norms between quantized and unquantized weights and report the best performing configuration. We also use that as initialization for LR-QAT. For weight-only quantization, we compare with GPTQ [18], AWQ [41], OmniQuant [55], and our implementation of PEQA [32], where we use symmetric weight quantization and following the same experimental setup and best RTN initialization, for a fair comparison. Similar to our method, we also apply PEQA only to linear layers in the attention blocks and keep token embeddings, final classification head and LayerNorm parameters frozen.
[…]
Method GPU memory, GB WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
Full-model QAT 71.9 5.74 69.29
LR-QAT 20.5 5.68 69.43
4.1 The impact of rank $r$
We investigate the effect of different values of rank $r$ of the auxiliary matrices $\bm{A}$ and $\bm{B}$ and present results in Figure 2. Increasing the rank from $1$ to $32$ leads to progressively slightly better performance, excluding one outlier. The fact that using $r>32$ doesn’t lead to further improvement in perplexity is likely because of the limited number of training steps we used for this experiment ($10^{3}$), and that more steps might be needed for the procedure to fully converge. Interestingly, a rank $r$ as small as 1 already performs really well. We hypothesize that this is the case because of the following. Even though $\operatorname{rank}\!\left(\bm{A}\bm{B}\right)=1$, by applying a low-rank approximation inside the rounding and clipping operators in (8), this can overall lead to a high-rank perturbation to the original weights $\bm{\Phi_{0}}$ (in the integer domain). Going forward, we use $r=32$ in all our experiments [^2] This amounts to only 1.2% of the total number of parameters for 7B LLaMA model..
@@ 4.1 The impact of rank $r$
Table 4: Weight-only quantization results for LLaMA-2/3 and Mistral models. We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better). Models marked ‘L2’/‘L3’, and ‘M’ denote LLaMA-2/3 and Mistral, respectively. Numbers marked in bold are SOTA or on par (within $0.05$). §Uses asymmetric weight quantization.
L2-7B L2-13B L3-8B M-7B L2-7B L2-13B L3-8B M-7B
FP16 5.47 4.88 6.14 5.25 70.47 73.18 74.22 75.69
W4pc RTN 6.14 5.21 7.53 5.91 68.88 71.73 72.19 73.44
GPTQ§ 5.83 5.13 - - - - - -
AWQ 6.15 5.12 - - - - - -
OmniQuant§ 5.74 5.02 - - - - - -
PEQA (our impl.) 5.71 5.03 7.51 5.56 69.23 72.51 72.79 73.73
LR-QAT (ours) 5.66 5.04 6.78 5.46 69.95 73.13 73.35 73.67
W3pc RTN 26.73 8.71 34.10 9.49 43.87 55.01 47.46 64.58
GPTQ§ 8.37 6.44 - - - - - -
AWQ 24.00 10.45 - - - - - -
OmniQuant§ 6.58 5.58 - - - - - -
PEQA (our impl.) 6.45 5.73 27.45 6.51 65.44 69.81 50.28 71.02
LR-QAT (ours) 6.13 5.54 8.12 6.03 67.66 71.22 70.46 71.87
W4g128 RTN 5.78 5.04 6.96 5.49 69.75 72.94 72.30 75.07
GPTQ§ 5.61 4.98 - - - - - -
AWQ 5.62 4.97 - - - - - -
OmniQuant§ 5.58 4.95 - - - - - -
PEQA (our impl.) 5.67 5.02 6.89 5.48 69.64 72.80 72.99 73.34
LR-QAT (ours) 5.59 4.97 6.62 5.36 69.88 72.89 73.72 75.18
W3g128 RTN 7.61 6.20 15.11 6.77 63.20 67.60 57.74 69.35
GPTQ§ 6.29 5.42 - - - - - -
AWQ 6.24 5.32 - - - - - -
OmniQuant§ 6.03 5.28 - - - - - -
PEQA (our impl.) 6.05 5.58 9.64 5.85 68.10 70.29 67.19 72.21
LR-QAT (ours) 5.99 5.53 7.74 5.81 67.98 70.54 70.48 71.44
@@ 4.4 Main results
We summarize our results in Table 4. As we can see, in almost most cases LR-QAT outperforms or is on par with prior weight-only quantization methods across various LLM families and quantization settings, including both per-channel and group-wise quantization.
@@
Table 5: Weight and activation quantization results for LLaMA-1 7B. We report WikiText-2 test set perplexity (lower is better) and zero-shot accuracy of 6 tasks (higher is better). Numbers marked in bold are SOTA. §Uses asymmetric weight quantization. *Uses a maximum sequence length of 1024 for evaluation.
# Bits (W-A-KV) Method WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
FP16 5.68 75.05 79.16 70.01 72.85 44.80 76.21 69.68
4-8-8 RTN 23.18 60.55 67.14 53.43 54.08 33.11 50.77 53.18
SmoothQuant 13.7* 71.00 76.00 66.00 67.40 42.80 67.80 65.17
LLM-QAT 11.2* 74.60 77.50 67.70 70.20 45.60 73.50 68.18
PEQA (our impl.) 5.89 74.86 78.24 70.01 70.12 42.83 75.14 68.53
LR-QAT (ours) 5.85 73.76 78.51 71.19 71.09 41.81 75.10 68.58
4-8-4 RTN 182.7 48.01 52.23 48.78 29.00 22.70 28.52 38.21
SmoothQuant 163.6* 54.70 55.40 51.50 43.90 27.70 38.90 45.35
LLM-QAT 11.6* 69.50 75.40 64.60 66.00 43.80 69.20 64.75
PEQA (our impl.) 6.15 72.97 77.80 67.72 67.13 40.27 73.35 66.54
LR-QAT (ours) 6.07 73.64 77.91 67.56 69.28 41.30 73.25 67.16
4-4-4 RTN 4.1e4 41.41 49.24 51.30 27.19 28.84 26.30 37.38
SmoothQuant 25.25 49.10 49.80 48.00 30.40 25.80 27.40 38.42
LLM-QAT - 61.30 51.50 51.90 27.90 23.90 31.10 41.27
LLM-QAT + SQ - 62.40 55.90 50.60 35.50 26.40 47.80 46.43
OS+ - 60.21 62.73 52.96 39.98 30.29 44.39 48.43
OmniQuant§ 11.26 63.51 66.15 53.43 45.20 31.14 56.44 52.65
PEQA (our impl.) 8.60 65.69 72.31 59.83 56.52 34.22 61.79 58.39
LR-QAT (ours) 8.47 67.16 71.76 59.59 58.42 34.73 62.34 59.00
We present our results for weight-activation quantization applied to LLaMA-1 7B in Table 5. LR-QAT demonstrates superior performance compared to all the PTQ and QAT baselines, consistently outperforming them across all the bitwidth settings. In addition to that, as we decrease the activation bitwidths, the improvement in model performance compared to prior work becomes more pronounced.
This indicates LR-QAT’s versatility, being readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic. Finally, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.

### v3

@@ Abstract
Large language models (LLMs) are omnipresent, however their practical deployment is challenging due to their ever increasing computational and memory demands. Quantization is one of the most effective ways to make them more compute and memory efficient. Quantization-aware training (QAT) methods, generally produce the best quantized performance, however it comes at the cost of potentially long training time and excessive memory usage, making it impractical when applying for LLMs. Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose LR-QAT – a lightweight and memory-efficient QAT algorithm for LLMs. LR-QAT employs several components to save memory without sacrificing predictive performance: (a) low-rank auxiliary weights that are aware of the quantization grid; (b) a downcasting operator using fixed-point or double-packed integers and (c) checkpointing. Unlike most related work, our method (i) is inference-efficient, leading to no additional overhead compared to traditional PTQ; (ii) can be seen as a general extended pretraining framework, meaning that the resulting model can still be utilized for any downstream task afterwards; (iii) can be applied across a wide range of quantization settings, such as different choices quantization granularity, activation quantization, and seamlessly combined with many PTQ techniques. We apply LR-QAT to LLaMA-1/2/3 and Mistral model families and validate its effectiveness on several downstream tasks. Our method outperforms common post-training quantization (PTQ) approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage. Specifically, we can train a 7B LLM on a single consumer grade GPU with 24GB of memory. Our source code is available at https://github.com/qualcomm-ai-research/LR-QAT.
@@ 1 Introduction
Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose Low-Rank Quantization-Aware Training (LR-QAT) – a lightweight memory-efficient and inference-efficient QAT algorithm for LLMs. LR-QAT reduces the memory requirements of training a 7B LLM from >98GB of GPU memory to <21GB without degrading the predictive performance compared to traditional full-model QAT, making it possible to train such models on a single consumer grade GPU. Unlike most related work that combines low-rank adaptation with quantization, our method is also inference-efficient. After the training is complete, the auxiliary matrices are naturally absorbed into the quantized weight tensor without loss of accuracy and no extra overhead at inference time. LR-QAT is positioned as a general extended pretraining method, as opposed to being strictly a fine-tuning method – the resulting model is a low-bit general pretrained LLM, that can still be utilized for any task afterwards. If needed, our resulting low-bit pretrained LLM can be fine-tuned on specific downstream tasks or used with multiple LoRA adapters for rapid switching between various tasks.
@@
We apply LR-QAT to LLaMA-1/2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks. Our method outperforms recent PTQ approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage. Finally, our method can be applied across a wide range of quantization settings, including per-channel or per-block weight quantization, activation quantization, and can be combined with most of other PTQ techniques.
@@ 2 Background and related work
Figure 1: Left: A schematic illustration of our proposed LR-QAT. $\mathbf{x}$ and $\mathbf{y}$ denote the input and the output of the linear layer. Right: Memory requirements for training with various QAT techniques on LLaMA-2 7B, assuming batch size 1, sequence length 1024, $r=32$, and BF16 compute data type. ‘Intermediate results’ refer to the results of some intermediate computations, e.g. after rounding/clipping in (3), which are saved in memory for the backward pass.
@@ 3 Method
Let’s recall how traditional QAT [17] works. Given a linear layer with a weight matrix $\bm{W}\in\mathbb{R}^{m\times k}$ and assuming $b$-bit symmetric uniform affine quantization, the quantization is simulated as follows:
@@
Note that both in the original LoRA paper [25] and in the related work like QLoRA [16], there is no need to compute the product $\bm{A}\bm{B}$ explicitly. Instead, those matrices are multiplied with the activations $\mathbf{x}$ as $\bm{A}\left(\bm{B}\mathbf{x}\right)$. However, we have to compute a product $\bm{A}\bm{B}$ in (5), and in the naïve implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage. To prevent this, we employ gradient checkpointing [10] on (5). In other words, we recompute the quantizer function in the backward pass, leading to a slight runtime overhead but avoiding significantly increased memory usage.
@@ 4 Experiments
We assess the effectiveness of LR-QAT by conducting experiments on LLaMA 7B [60], LLaMA-2 7B/13B [61], LLaMA-3 8B [1], and Mistral-0.1 7B [31]. We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\!\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$. We then compare our method in terms of accuracy to standard full-model QAT, other baselines, and the related work. All detailed hyperparameters of our experiments are in Appendix B.
Quantization
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 / INT2 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We use symmetric quantization, except the INT2 case, where we use asymmetric quantization (1), for a fair comparison with related work. We quantize all linear layers, except the classification head. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [55], we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
Datasets and training
We apply our method to all linear layers in the attention blocks, both in self-attention and in the feed-forward network. We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$ and keep embedding layers, final classification head and RMSNorm parameters frozen. In the case of asymmetric weight quantization, a zero offset $\mathbf{z}$ is set during range estimation phase and kept frozen throughout training ($\mathbf{z}=\mathbf{z_{0}}$).
We train on a small subset of SlimPajama [57], which is an open-source dataset similar to the original one used for pretraining LLaMA models. In all experiments we train using batch size $32$ and a maximum sequence length of $1024$. For all weight-only and weight-activation quantization results, we train for $10^{4}$ steps. For ablation studies in Sections 4.1 and 4.2 we use shorter training of $10^{3}$ steps. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
Evaluation
Following the previous work [18, 65, 55, 44], we evaluate quantized models by reporting the perplexity of language generation on WikiText-2 [48], using a sequence length of $2048$. We also report zero-shot accuracy on a set of common sense reasoning tasks including BoolQ [12], PIQA [5], Winogrande [54], ARC [13], and HellaSwag [69]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19]. Specifically, we use lm_eval v0.4.2 and report acc_norm for tasks where it’s available (PIQA, ARC-e, ARC-c, HellaSwag) and otherwise acc (BoolQ and Winogrande).
Baselines
We compare with round-to-nearest quantization (RTN), where we set the ranges based on minimizing the $L^{p}$-norms between quantized and unquantized weights and report the best performing configuration. We also use that as initialization for LR-QAT. For weight-only quantization, we compare with GPTQ [18], AWQ [41], and OmniQuant [55]. We also compare with our implementation of PEQA [32] and full-model QAT (LSQ) [17], where we follow the same experimental setup as for our method, together with the best RTN initialization, for a fair comparison.
4.1 The impact of rank $r$
@@
$\varphi\!\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. WikiText-2 $\downarrow$ Zero-shot acc. $\uparrow$
W4 pc W3 pc W4 pc W3 pc
FP32 FP32 LoRA 5.69 6.21 69.28 66.62
FP16 FP32 LoRA $+$0.00 $+$0.01 --0.13 --0.01
BF16 FP32 LoRA --0.01 $+$0.01 $+$0.11 $+$0.45
Q4.4 / Q3.5 FP32 LoRA --0.01 $+$0.01 $+$0.16 $+$0.31
Q4.4 / Q3.5 BF16 LoRA --0.01 $+$0.01 $+$0.15 $+$0.31
INT-4 / INT-3 FP32 LoRA $+$0.02 $+$20.5 --0.04 --22.8
INT-4 / INT-3 FP32 LoftQ ($T=1$) $+$0.28 $+$0.18 --0.67 $+$0.26
INT-4 / INT-3 FP32 LoftQ ($T=64$) $+$0.40 $+$1.37 --1.40 --2.01
@@
We investigate the effect of different values of rank $r$ of the auxiliary matrices $\bm{A}$ and $\bm{B}$ and present results in Figure 2. Increasing the rank from $1$ to $32$ leads to progressively slightly better performance, excluding one outlier. The fact that using $r>32$ doesn’t lead to further improvement in perplexity is likely because of the limited number of training steps we used for this experiment ($10^{3}$), and that more steps might be needed for the procedure to fully converge. Interestingly, a rank $r$ as small as 1 already performs really well. We hypothesize that this is the case because of the following. Even though $\operatorname{rank}\!\left(\bm{A}\bm{B}\right)=1$, by applying a low-rank approximation inside the rounding and clipping operators in (8), this can overall lead to a high-rank perturbation to the original weights $\bm{\Phi_{0}}$ (in the integer domain). Finally, for all ranks we observe only a small standard deviation between $0.005$ and $0.008$ ppl., indicating the robustness of LR-QAT to a random initialization of $\bm{B}$. Going forward, we use $r=32$ in all our experiments [^2] This amounts to only 1.2% of the total number of parameters for 7B LLaMA model..
@@ 4.3 Comparison with full-model QAT
Method GPU mem., Time/100 steps, WikiText-2 ppl. $\downarrow$ Zero-shot acc. $\uparrow$
GB sec W4 pc W3 pc W4 pc W3 pc
Full-model QAT (LSQ) 62.2 (98.5) 3248±${}^{\scriptsize\pm{}\text{7}}$ 5.77±${}^{\scriptsize\pm{}\text{0.02}}$ 6.14±${}^{\scriptsize\pm{}\text{0.01}}$ 68.96±${}^{\scriptsize\pm{}\text{0.29}}$ 67.14±${}^{\scriptsize\pm{}\text{0.13}}$
LR-QAT (ours) 20.5 1522±${}^{\scriptsize\pm{}\text{5}}$ 5.66±${}^{\scriptsize\pm{}\text{0.00}}$ 6.13±${}^{\scriptsize\pm{}\text{0.02}}$ 69.72±${}^{\scriptsize\pm{}\text{0.32}}$ 67.70±${}^{\scriptsize\pm{}\text{0.25}}$
[…]
As we can see in Table 3, training with our method leads to on par or better predictive performance at a significantly lower memory usage and runtime compared to full-model QAT. We include results for other models and bitwidths in Table 4. A more detailed runtime comparison can also be found in Table A8.
@@ 4.4 Main results
Table 4: Weight-only quantization results for LLaMA and Mistral models. We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better). Models marked ‘L1’/‘L2’/‘L3’, and ‘M’ denote LLaMA-1/2/3 and Mistral, respectively. Numbers marked in bold are SOTA or on par (within $0.05$). §Uses asymmetric weight quantization. ✕ denotes out of memory.
L1-7B L2-7B L2-13B L3-8B M-7B L1-7B L2-7B L2-13B L3-8B M-7B
FP16 5.68 5.47 4.88 6.14 5.25 69.68 70.47 73.18 74.22 75.69
W4 pc RTN 6.33 6.14 5.21 7.53 5.91 68.51 68.88 71.73 72.19 73.44
GPTQ§ 6.13 5.83 5.13 - - 64.95 - - - -
AWQ 6.08 6.15 5.12 - - - - - - -
OmniQuant§ 5.86 5.74 5.02 - - - - - - -
LSQ (our impl.) 5.94 5.77 6.87 5.73 68.37 68.96 73.28 72.88
PEQA (our impl.) 5.86 5.71 5.03 7.51 5.56 68.49 69.23 72.51 72.79 73.73
LR-QAT (ours) 5.84 5.66 5.03 6.78 5.46 68.54 69.72 73.19 73.84 74.44
W4 g128 RTN 6.05 5.78 5.04 6.96 5.49 68.93 69.75 72.94 72.30 75.07
GPTQ§ 5.85 5.61 4.98 - - - - - - -
AWQ 5.81 5.62 4.97 - - - - - - -
OmniQuant§ 5.77 5.58 4.95 - - - - - - -
LSQ (our impl.) 5.76 5.61 6.58 5.67 69.17 69.68 73.31 72.90
PEQA (our impl.) 5.75 5.67 5.02 6.89 5.48 69.19 69.64 72.80 72.99 73.34
LR-QAT (ours) 5.75 5.59 4.97 6.57 5.37 69.15 69.88 72.91 73.66 75.28
W3 pc RTN 12.88 26.73 8.71 34.10 9.49 54.66 43.87 55.01 47.46 64.58
GPTQ§ 8.06 8.37 6.44 - - - - - - -
AWQ 11.88 24.00 10.45 - - - - - - -
OmniQuant§ 6.49 6.58 5.58 - - - - - - -
LSQ (our impl.) 6.29 6.14 8.14 6.06 66.29 67.14 69.58 71.61
PEQA (our impl.) 6.56 6.45 5.73 26.20 6.51 65.75 65.44 69.81 51.05 71.02
LR-QAT (ours) 6.27 6.13 5.54 8.12 6.03 66.60 67.70 71.22 70.46 71.87
W3 g128 RTN 7.96 7.61 6.20 15.11 6.77 63.50 63.20 67.60 57.74 69.35
GPTQ§ 6.55 6.29 5.42 - - - - - - -
AWQ 6.46 6.24 5.32 - - - - - - -
OmniQuant§ 6.15 6.03 5.28 - - - - - - -
LSQ (our impl.) 6.20 6.02 8.08 5.90 66.53 68.36 70.11 71.96
PEQA (our impl.) 6.22 6.05 5.58 9.64 5.85 66.66 68.10 70.29 67.19 72.21
LR-QAT (ours) 6.17 5.98 5.32 7.74 5.80 66.81 68.62 71.51 70.48 72.41
W2 pc RTN§ 4.9e3 5.2e3 5.2e3 6.4e4 6.8e3 37.92 36.52 36.27 36.80 36.59
GPTQ§ 2.1e3 7.7e3 2.1e3 - - - - - - -
OmniQuant§ 15.47 37.37 17.21 - - - - - - -
PEQA (our impl.)§ 8.24 9.34 7.51 14.47 8.08 59.83 57.40 62.29 57.72 64.48
LR-QAT (ours)§ 7.99 8.51 7.16 12.52 7.98 61.77 60.03 65.28 58.49 65.11
W2 g128 RTN§ 708 2.5e3 115.6 1.4e4 369 39.74 37.94 41.12 36.97 41.30
GPTQ§ 44.01 36.77 28.14 - - - - - - -
AWQ 2.6e5 2.2e5 1.2e5 - - - - - - -
OmniQuant§ 9.72 11.06 8.26 - - - - - - -
PEQA (our impl.)§ 8.12 7.87 6.83 12.39 8.05 60.69 60.74 64.74 57.81 64.65
LR-QAT (ours)§ 7.86 7.62 6.57 11.09 7.92 61.60 61.70 66.75 60.46 65.27
@@
We summarize our results in Table 4. As we can see, in almost most cases LR-QAT outperforms or is on par with prior weight-only quantization methods across various LLM families and quantization settings, including both per-channel and group-wise quantization. In the case of extremely low-bitwidth regime, our method consistently and significantly outperforms related work, across all settings.
@@
Table 5: Weight and activation quantization results for LLaMA-1/2 (denoted ‘L1’/‘L2’, respectively). We report WikiText-2 test set perplexity and zero-shot accuracy of 6 tasks. Numbers marked in bold are SOTA. §Uses asymmetric weight quantization. *Uses a maximum sequence length of 1024 for evaluation.
# Bits (W-A-KV) Method WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
L1-7B L2-7B L2-13B L1-7B L2-7B L2-13B
FP16 5.68 5.47 4.88 69.68 70.47 73.18
4-8-8 RTN 6.88 6.17 5.23 65.83 68.55 71.54
SmoothQuant 13.7* - - 65.17 - -
LLM-QAT 11.2* - - 68.18 - -
PEQA (our impl.) 5.89 5.72 5.08 68.53 69.11 72.49
LR-QAT (ours) 5.85 5.67 5.04 68.58 69.32 73.18
4-8-4 RTN 7.66 6.85 5.78 62.78 65.16 67.06
SmoothQuant 163.6* - - 45.35 - -
LLM-QAT 11.6* - - 64.75 - -
PEQA (our impl.) 6.15 6.03 5.27 66.54 67.56 71.35
LR-QAT (ours) 6.07 5.90 5.24 67.16 69.84 71.65
4-4-4 RTN 17.75 18.98 11.37 49.60 51.75 55.07
SmoothQuant 25.25 83.12 35.88 38.42 - -
LLM-QAT - - - 41.27 - -
LLM-QAT + SQ - - - 46.43 - -
Outlier Suppression+ - - - 48.43 - -
OmniQuant§ 11.26 14.26 12.30 52.65 - -
PEQA (our impl.) 8.60 8.72 7.23 58.39 57.93 62.26
LR-QAT (ours) 8.47 8.46 7.15 59.00 58.98 62.65
We present our results for weight-activation quantization applied to LLaMA-1/2 models in Table 5. LR-QAT demonstrates superior performance compared to all the PTQ and QAT baselines, consistently outperforming them across all model families and the bitwidth settings. In addition to that, as we decrease the activation bitwidths, the improvement in model performance compared to prior work becomes more pronounced.
This indicates LR-QAT’s versatility, being readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic. Further, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.